OpenStudy (anonymous):

i need help on a question that i have been stuck on forever, please help! i'll give you meddles! 2x + 3y = 1,200 Change the equation into slope-intercept form. Identify the slope and y-intercept of the equation. & I know this is a pain but please dont just tell me the answer can you show me how to work it out, please?

OpenStudy (anonymous):

slope intercept form is: y = mx + b We need to isolate y. 2x + 3y = 1200 3y = -2x + 1200 y = -2/3x + 400

OpenStudy (anonymous):

y = mx + b m is the slope b is the y intercept So... m = -2/3 b = 400 y = (-2/3)x + (400)

OpenStudy (anonymous):

I get to that point then i don't know what to do what the fraction..o.0

OpenStudy (anonymous):

This is what i got: 2x+3y=1200 -2x -2x ---------------- 3y=-2x+1200 3 3 3 ----------------- y=-2/3x+400 Now i don't know where to go from here...can you please help me? @jdoe0001 @jcpd910

OpenStudy (anonymous):

Ah, you don't have to do anything! :D The slope is Rise/Run, so your slope is -2/3 That means you go down 2, over 3, and plot that point. :D

OpenStudy (anonymous):

what is the y-intercept?

OpenStudy (anonymous):

y = mx + b, where b is the y intercept. 400 is b in this equation.

OpenStudy (anonymous):

Okay thank you! can you help me on my other question please? its similar. :D

OpenStudy (jdoe0001):

\(\bf 2x + 3y = 1,200 \qquad {\color{brown}{ -2x}} \\ \quad \\ \cancel{ {\color{brown}{ -2x}}+2x } + 3y = {\color{brown}{ -2x}}+1,200\qquad {\color{brown}{ \div 3}} \\ \quad \\ \cfrac{\cancel{ 3 }y}{\cancel{ {\color{brown}{ 3}} }}=\cfrac{-2x+1200}{{\color{brown}{ 3}}}\implies y=\cfrac{-2x}{3}+\cancel{ \cfrac{1200}{3} } \\ \quad \\ y={\color{purple}{ -\cfrac{2}{3}}}x{\color{blue}{ +400}}\qquad {\color{purple}{ slope}}\ and \ {\color{blue}{ y-intercept}}\)

OpenStudy (anonymous):

Unfortunately I have to go, but jdoe can help. :)

OpenStudy (anonymous):

Okay, thank you again!